OpenStudy (anonymous):

Use the definition of limit and properties of limit to show that function is continuous at the given number f(x)=x2+√7-x , a=4

OpenStudy (anonymous):

I'm not sure whether I've read the function in the right way, but I think you mean \(f(x) =x^2 + \sqrt{7-x}\) and not \(f(x) = 2x + \sqrt{7} -x\) If it's correct, you can see that \(f(x)\) is composed of 2 terms having x as a variable. So you can view each of them as an independent function: \(x^2\) and \(\sqrt{7-x}\). This will help us to see how to apply the properties of the limits. By definition, \(f(x)\) will be continuous at \(a=4\) if \(\lim_{x \to a} f(x)\) exists and it's equal to \(f(a)\). On the left-hand side we're going to prove that in the faster way, by plugging 4 in \(f(x)\) and on the left we're going to do the same but by using the limit properties. Let use 4 instead of a: \[\large{\lim_{x \to 4} (x^2 + \sqrt{7-x}) = \lim_{x \to 4} (x^2 + \sqrt{7-x})}\] \[\large{4^2 + \sqrt{7-4} = \lim_{x \to 4} (x^2) + \lim_{x \to 4}(\sqrt{7-x})\hspace{35pt} (1)}\] \[\large{16 + \sqrt{3} = \left(\lim_{x \to 4} x\right)^2 + \sqrt{\lim_{x \to 4}(7-x)}\hspace{35pt} (2)}\] \[\large{16 + \sqrt{3} = 16 + \sqrt{7-4}}\hspace{35pt} (3)\] \[\large{16 + \sqrt{3} = 16 + \sqrt{3}}\hspace{35pt} (4)\] \((1)\): Take the limit as if it were a sum of two functions. \((2)\): The property for functions rise to some number: the first to the 2nd power and the second to the \(\frac{1}{2}\) power. \((3)\): Just arithmetic. \((4)\): We're done.